# Correcting the Mythos of KL-Regularization: Direct Alignment without Overoptimization via $\chi^{2}$-Preference Optimization

### v2

@@ G.3 Proofs for Supporting Lemmas
Lemma G.4 (three point lemma).
@@
Substitute the definition into Lemma G.4 and we can prove the lemma. ∎
From Lemma G.4, we can rewrite (7) as follows:
@@
Proof of Lemma G.3 Let $\widehat{L}(\pi,\pi^{\prime},\pi^{\prime\prime})$ denote the empirical squared loss:
@@
Then from Lemma G.2, we know with probability at least $1-\delta^{\prime}$ that
@@
The section contains the proofs for the main guarantee $\chi^{2}$-RLHF in Appendix C (Theorem C.1). We first prove two results, Theorem H.1 and Corollary H.1, which correspond to exact (i.e., including precise constants) versions of the two statements in Theorem C.1. We also analyze $\chi^{2}$-RLHF with $\eta=0$ in Corollary H.2.
Throughout this section, we make use of the following $\eta$-smoothed version of the $L_{1}$ concentrability (Definition 3.1) coefficient:
@@
Theorem H.1 (General regret bound for Algorithm 3).
Suppose Assumption C.1 and Assumption C.2 hold for parameters $\beta>0$ and $\eta\in\big{[}0,\frac{\beta}{8R_{\mathsf{max}}}\big{]}$. Then with probability at least $1-\delta$, the policy $\widehat{\pi}$ produced by $\chi^{2}$-RLHF (Algorithm 3) satisfies
@@
The following results are immediate consequences of Theorem H.1.
Corollary H.1 (Smoothed $\chi^{2}$-regularization).
Given $\pi^{\star}$, let $\eta=\frac{\beta}{8R_{\mathsf{max}}}$ and $\beta=2\sqrt{\frac{32R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\log(3|\mathcal{R}|/\delta)}{n\mathcal{C}^{\pi^{\star}}}}$. Then under the preconditions of Theorem H.1, with probability at least $1-\delta$, the policy $\widehat{\pi}$ produced by $\chi^{2}$-RLHF (Algorithm 3) satisfies
@@
Corollary H.2 (Non-smoothed $\chi^{2}$-regularization).
Given $\pi^{\star}$, let $\eta=0$ and $\beta=2\sqrt{\frac{32R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\log(3|\mathcal{R}|/\delta)}{n\mathcal{C}^{\pi^{\star}}}}$. Then under the preconditions of Theorem H.1, with probability at least $1-\delta$, the policy $\widehat{\pi}$ produced by $\chi^{2}$-RLHF (Algorithm 3) satisfies
@@
Proof of Theorem H.1 The proof follows largely the same lines of analyses as the proof of Theorem E.1. One difference is that in Algorithm 3, we approximate the RLHF objective using contexts are sampled from $\mathcal{D}_{\mathsf{x}}$, so we require additional concentration arguments to show that the empirical objective approximates its population counterpart.
@@
Next, we state the following off- and on-policy reward estimation error bounds for the reward model ${\widehat{r}}$, analogous to Lemma E.1 and Lemma E.3 for $\chi$PO.
Lemma H.2.
Suppose Assumption C.1 holds. Then with probability at least $1-\delta$, the reward model ${\widehat{r}}$ learned in Eq. 36 satisfies
@@
Lemma H.3.
Under the event in Lemma H.2, we have that for all $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$,
@@
where $\varepsilon_{\mathrm{stat}}^{2}$ is defined in Lemma H.2.
@@
Equipped with these concentration and estimation error bounds, we now bound the regret of Algorithm 3 using a pessimism-based analysis similar to the proof of Theorem E.1. Condition on the events in Eq. 56, Lemma H.1, and Lemma H.2, which hold together with probability at least $1-3\delta^{\prime}$. We decompose the regret of $\widehat{\pi}$ using $\widehat{J}_{\beta,\eta}^{{{\widehat{r}}}}$, then leverage the inequalities in Eq. 57 and Eq. 58:
@@
In the last line above, we have introduced the notation $\Delta^{\star}(x,a,b)=r^{\star}(x,a)-r^{\star}(x,b)$ and $\widehat{\Delta}(x,a,b)={\widehat{r}}(x,a)-{\widehat{r}}(x,b)$, and centered the returns. Next, applying Lemma H.3 to bound the reward estimation error above, we have
@@
Proof of Corollary H.1 When $\eta=\frac{\beta}{8R_{\mathsf{max}}}$, Theorem H.1 states that
@@
Proof of Corollary H.2 When $\eta=0$, Theorem H.1 states that
@@
Proof of Lemma H.2 We use similar reasoning and notation to the proof of Lemma E.1. Since $r^{\star}\in\mathcal{R}$ under Assumption C.1, Lemma D.1 guarantees that with probability at least $1-\delta$ we have
@@
Since $|r(x,a)-r(x,b)|\leq R_{\mathsf{max}}$ for all $r\in\mathcal{R}$ under Assumption C.1, we then apply Lemma E.5 with $R=V=R_{\mathsf{max}}$.
@@
Proof of Lemma H.3 Abbreviate $\Delta^{\star}(x,a,b)=r^{\star}(x,a)-r^{\star}(x,b)$, and $\widehat{\Delta}(x,a,b)={\widehat{r}}(x,a)-{\widehat{r}}(x,b)$. For a pair of policies $\pi,\pi^{\prime}$ and $p\geq 1$, we define the norm $\|\cdot\|_{p,\pi\times\pi^{\prime}}\vcentcolon={}\left(\operatorname{\mathbb{E}}_{\rho,a\sim\pi,b\sim\pi^{\prime}}[|\cdot|^{p}]\right)^{1/p}$, so that $\operatorname{\mathbb{E}}_{\pi,\pi_{\mathsf{ref}}}\left[\left\lvert\Delta^{\star}(x,a,b)-\widehat{\Delta}(x,a,b)\right\rvert\right]=\left\|\Delta^{\star}-\widehat{\Delta}\right\|_{1,\pi\times\pi_{\mathsf{ref}}}$. Then via Cauchy-Schwarz,

### v3

@@ H.3 Proofs for Supporting Lemmas
Lemma H.4 (three point lemma).
@@
Substitute the definition into Lemma H.4 and we can prove the lemma. ∎
From Lemma H.4, we can rewrite (7) as follows:
@@
Proof of Lemma H.3 Let $\widehat{L}(\pi,\pi^{\prime},\pi^{\prime\prime})$ denote the empirical squared loss:
@@
Then from Lemma H.2, we know with probability at least $1-\delta^{\prime}$ that
@@
The section contains the proofs for the main guarantee $\chi^{2}$-RLHF in Appendix B (Theorem B.1). We first prove two results, Theorem I.1 and Corollary I.1, which correspond to exact (i.e., including precise constants) versions of the two statements in Theorem B.1. We also analyze $\chi^{2}$-RLHF with $\eta=0$ in Corollary I.2.
Throughout this section, we make use of the following $\eta$-smoothed version of the $L_{1}$ concentrability coefficient:
@@
Theorem I.1 (General regret bound for Algorithm 3).
Suppose Assumption B.1 and Assumption B.2 hold for parameters $\beta>0$ and $\eta\in\big{[}0,\frac{\beta}{8R_{\mathsf{max}}}\big{]}$. Then with probability at least $1-\delta$, the policy $\widehat{\pi}$ produced by $\chi^{2}$-RLHF (Algorithm 3) satisfies
@@
The following results are immediate consequences of Theorem I.1.
Corollary I.1 (Smoothed $\chi^{2}$-regularization).
Given $\pi^{\star}$, let $\eta=\frac{\beta}{8R_{\mathsf{max}}}$ and $\beta=2\sqrt{\frac{32R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\log(3|\mathcal{R}|/\delta)}{n\mathcal{C}^{\pi^{\star}}}}$. Then under the preconditions of Theorem I.1, with probability at least $1-\delta$, the policy $\widehat{\pi}$ produced by $\chi^{2}$-RLHF (Algorithm 3) satisfies
@@
Corollary I.2 (Non-smoothed $\chi^{2}$-regularization).
Given $\pi^{\star}$, let $\eta=0$ and $\beta=2\sqrt{\frac{32R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\log(3|\mathcal{R}|/\delta)}{n\mathcal{C}^{\pi^{\star}}}}$. Then under the preconditions of Theorem I.1, with probability at least $1-\delta$, the policy $\widehat{\pi}$ produced by $\chi^{2}$-RLHF (Algorithm 3) satisfies
@@
Proof of Theorem I.1 The proof follows largely the same lines of analyses as the proof of Theorem F.1. One difference is that in Algorithm 3, we approximate the RLHF objective using contexts are sampled from $\mathcal{D}_{\mathsf{x}}$, so we require additional concentration arguments to show that the empirical objective approximates its population counterpart.
@@
Next, we state the following off- and on-policy reward estimation error bounds for the reward model ${\widehat{r}}$, analogous to Lemma F.1 and Lemma F.3 for $\chi$PO.
Lemma I.2.
Suppose Assumption B.1 holds. Then with probability at least $1-\delta$, the reward model ${\widehat{r}}$ learned in Eq. 33 satisfies
@@
Lemma I.3.
Under the event in Lemma I.2, we have that for all $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$,
@@
where $\varepsilon_{\mathrm{stat}}^{2}$ is defined in Lemma I.2.
@@
Equipped with these concentration and estimation error bounds, we now bound the regret of Algorithm 3 using a pessimism-based analysis similar to the proof of Theorem F.1. Condition on the events in Eq. 56, Lemma I.1, and Lemma I.2, which hold together with probability at least $1-3\delta^{\prime}$. We decompose the regret of $\widehat{\pi}$ using $\widehat{J}_{\beta,\eta}^{{{\widehat{r}}}}$, then leverage the inequalities in Eq. 57 and Eq. 58:
@@
In the last line above, we have introduced the notation $\Delta^{\star}(x,a,b)=r^{\star}(x,a)-r^{\star}(x,b)$ and $\widehat{\Delta}(x,a,b)={\widehat{r}}(x,a)-{\widehat{r}}(x,b)$, and centered the returns. Next, applying Lemma I.3 to bound the reward estimation error above, we have
@@
Proof of Corollary I.1 When $\eta=\frac{\beta}{8R_{\mathsf{max}}}$, Theorem I.1 states that
@@
Proof of Corollary I.2 When $\eta=0$, Theorem I.1 states that
@@
Proof of Lemma I.2 We use similar reasoning and notation to the proof of Lemma F.1. Since $r^{\star}\in\mathcal{R}$ under Assumption B.1, Lemma E.1 guarantees that with probability at least $1-\delta$ we have
@@
Since $|r(x,a)-r(x,b)|\leq R_{\mathsf{max}}$ for all $r\in\mathcal{R}$ under Assumption B.1, we then apply Lemma F.5 with $R=V=R_{\mathsf{max}}$.
@@
Proof of Lemma I.3 Abbreviate $\Delta^{\star}(x,a,b)=r^{\star}(x,a)-r^{\star}(x,b)$, and $\widehat{\Delta}(x,a,b)={\widehat{r}}(x,a)-{\widehat{r}}(x,b)$. For a pair of policies $\pi,\pi^{\prime}$ and $p\geq 1$, we define the norm $\|\cdot\|_{p,\pi\times\pi^{\prime}}\vcentcolon={}\left(\operatorname{\mathbb{E}}_{\rho,a\sim\pi,b\sim\pi^{\prime}}[|\cdot|^{p}]\right)^{1/p}$, so that $\operatorname{\mathbb{E}}_{\pi,\pi_{\mathsf{ref}}}\left[\left\lvert\Delta^{\star}(x,a,b)-\widehat{\Delta}(x,a,b)\right\rvert\right]=\left\|\Delta^{\star}-\widehat{\Delta}\right\|_{1,\pi\times\pi_{\mathsf{ref}}}$. Then via Cauchy-Schwarz,